A tuna fish can (4.0 cm radius) rolls at 2.5 m/s on a level surface and encounters a 45'' incline. How far up the incline does it roll?
Answer #1

Usinh conservation of energy

Potential energy possessed = rotational KE + translational kE

mg h = 0.5 I w^2 + 0.5 m v^2

m *9.8* (d sin 45) = 0.5* (0.5 m r^2) ( v/r)^2 + 0.5 m v^2

9.8* d sin 45 = 0.25 * 2.5^2 + 0.5* 2.5^2

d = 0.676 m

=========

Comment in case any doubt, will reply for sure.. Goodluck